OpenStudy (lilmike234):

if logx=-2 what is x

OpenStudy (rhr12):

-2/log

OpenStudy (e.mccormick):

if \(y=b^x\) then \(\log _b(y)=x\). You can use that relationship to move things around. that is one way to go at it.

OpenStudy (rhr12):

-2/log=x

OpenStudy (lilmike234):

So the answer is 0.01 correct ? @e.mccormick

OpenStudy (e.mccormick):

Yes.